OpenStudy (anonymous):

find y' by implicit differentiation. 7/x+7/y=4

OpenStudy (luigi0210):

Any attempts ?

zepdrix (zepdrix):

It's so hard to read that with all the fraction bars. Is this the function? \[\Large\rm \frac{7}{x}+\frac{7}{y}=4\] You can write it like this if it's easier to apply your power rules,\[\Large\rm 7x^{-1}+7y^{-1}=4\]

zepdrix (zepdrix):

What part are you getting stuck on? :U

OpenStudy (anonymous):

am i suppose to put this in implicit form before doing the power rule?

OpenStudy (dumbcow):

\[\rightarrow -7x^{-2} -7y^{-2}(y') = 0\] that is the differentiation now solve for y' \[y' = \frac{7x^{-2}}{-7y^{-2}} = -\frac{y^2}{x^{2}}\]